OpenStudy (anonymous):

The function for the cost of materials to make a shirt is f(x) = five sixths x + 5, where x is the number of shirts. The function for the selling price of those shirts is g(f(x)), where g(x) = 5x + 6. Find the selling price of 18 shirts help??!!!!!

jimthompson5910 (jim_thompson5910):

what is f(18)

OpenStudy (anonymous):

i'm lost?

jimthompson5910 (jim_thompson5910):

plug x = 18 into the function f(x)

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

96?

jimthompson5910 (jim_thompson5910):

f(x) = (5/6)*x + 5 f(18) = (5/6)*18 + 5 f(18) = ??

OpenStudy (anonymous):

20

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now plug that into g(x)

OpenStudy (anonymous):

plug 20 into g(x)??

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

106, got it,. thank you!

jimthompson5910 (jim_thompson5910):

yw